Shell Programming—an Introduction

Copyright Conditions: GNU FDL (see http://www.gnu.org/licenses/fdl.html  )

Nick Urbanik

                                 nicku@nicku.org

A computing department

Aim

After successfully working through this exercise, You will: write simple shell scripts using for, if, while, case, getopts statements; write shell script functions, and be able to handle parameters; understand basic regular expressions, and be able to create your own regular expressions; understand how to execute and debug these scripts; understand some simple shell scripts written by others.

Why Shell Scripting?

Basic startup, shutdown of Linux, Unix systems uses large number of shell scripts

understanding shell scripting important to

understand and perhaps modify behaviour of system

Very high level: powerful script can be very short

Can build, test script incrementally

Useful on the command line: “one liners”

Where to get more information

the Libarary has two copies of the book, Learning the Bash Shell, second edition, by Cameron Newham & Bill Rosenblatt, O’Reilly, 1998.

There is a free on-line book about shell programming

at: http://tldp.org/LDP/abs/html/index.html
and http://tldp.org/LDP/abs/abs-guide.pdf  .

It has hundreds of pages, and is packed with examples.

The handy reference to shell programming is:

$ pinfo bash

or

$ man bash

The Shell is an Interpreter

Some languages are compiled: C, C++, Java,…

Some languages are interpreted: Java bytecode, Shell

Shell is an interpreter: kernel does not run shell program directly:

kernel runs the shell program /bin/sh with script file name as a parameter the kernel cannot execute the shell script directly, as it can a binary executable file that results from compiling a C program

The Shebang

Linux kernel starts shell script kernel reads first two characters of each executable

If first 2 chars are “#!” then

kernel executes the name that follows, with the file name of the script as a parameter

Example: a file called find.sh has this as the first line:

#! /bin/sh

then kernel executes this:

/bin/sh find.sh

Making the script executable

To easily execute a script, it should: be on the PATH have execute permission.

How to do each of these?

Red Hat Linux by default, includes the directory ∼/bin on the PATH, so create this directory, and put your scripts there:

$ mkdir ∼/bin

If your script is called script, then this command will make it executable:

$ chmod +x script

Special Characters

Many characters are special to the shell, and have a particular meaning to the shell.

Character Meaning  See slide
∼ Home directory § 7
`  Command substitution (depracated) § 24
#  Comment
$  Variable expression  § 15
&  Background Job  ?? on page ??
*  File name matching wildcard  ?? on page ??
|  Pipe  ?? on page ??

OSSI — Shell Programming — ver. 1.9 – p. 8/85

Special Characters—continued: 2

Character Meaning  See slide
(  Start subshell  § 45, 17, 39
)  End subshell  § 45, 17, 39
[  Start character set file name matching  ?? on page ??
]  End character set file name matching  ?? on page ??
{  Start command block  § 39
;  Command separator  § 40
\  Quote next character  § 23
'  Strong quote  § 23
"  Weak quote  § 23

Special Characters—continued: 3

Character Meaning  See slide
<  Input redirect  ?? on page ??
>  Output redirect  ?? on page ??
/  Pathname directory separator
?  Single-character match in filenames  ?? on page ??
!  Pipline logical NOT  § 28

Note that references to pages in the tables above refer to the modules in the workshop notes

Quoting

Sometimes you want to use a special character literally; i.e., without its special meaning.

Called quoting

Suppose you want to print the string: 2 * 3 > 5 is

a valid inequality?

If you did this:

$ echo 2 * 3 > 5 is a valid inequality

the new file ‘5’ is created, containing the character ‘2’, then the names of all the files in the current directory,

then the string “3 is a valid inequality”.

Quoting—2

To make it work, you need to protect the special characters ‘*’ and ‘>’ from the shell by quoting them. There are three methods of quoting:

Using double quotes (“weak quotes”)

Using single quotes (“strong quotes”)

Using a backslash in front of each special character you want to quote

This example shows all three:

$ echo "2 * 3 > 5 is a valid inequality"
$ echo '2 * 3 > 5 is a valid inequality'
$ echo 2 \* 3 \> 5 is a valid inequality

Quoting—When to use it?

Use quoting when you want to pass special characters to another program.

Examples of programs that often use special characters:

   find, locate, grep, expr, sed and echo

Here are examples where quoting is required for the program to work properly:

$ find . -name \*.jpg
$ locate '/usr/bin/c*'
$ grep 'main.*(' *.c
$ i=`expr i \* 5`

True and False

Shell programs depend on executing external programs When any external program execution is successful, the exit status is zero, 0 An error results in a non-zero error code To match this, in shell programming:

The value 0 is true

any non-zero value is false This is opposite from other programming languages

Variables—1

Variables not declared; they just appear when assigned to Assignment:

no dollar sign no space around equals sign examples:

$ x=10 # correct $ x = 10 # wrong: try to execute program called ‘‘x’’

Read value of variable:

put a ‘$’ in front of variable name example:

$ echo "The value of x is $x"

Variables—Assignments

You can put multiple assignments on one line:

i=0 j=10 k=100

You can set a variable temporarily while executing a program:

$ echo $EDITOR
emacsclient
$ EDITOR=gedit crontab -e
$ echo $EDITOR
emacsclient

Variables—Local to Script

Variables disappear after a script finishes

Variables created in a sub shell disappear parent shell cannot read variables in a sub shell example:

   $ cat variables
   #! /bin/sh
   echo $HOME
   HOME=happy
   echo $HOME
   $ ./variables
   /home/nicku
   happy
   $ echo $HOME
   /home/nicku

Variables—Unsetting Them

You can make a variable hold the null string by assigning it to nothing, but it does not disappear totally:

$ VAR=
$ set | grep '^VAR'
VAR=

You can make it disappear totaly using unset:

$ unset VAR
$ set | grep '^VAR'

Command-line Parameters

Command-line parameters are called $0, $1, $2, … Example: when call a shell script called

“shell-script” like this:
$ shell-script param1 param2 param3 param4

variable value

$0       shell-script
$1       param1
$2       param2
$3       param3
$4       param4

$#  number of parameters to the program, e.g., 4

Note: these variables are read-only.

OSSI — Shell Programming — ver. 1.9 – p. 19/85

Special Built-in Variables

Both $@ and $* are a list of all the parameters.

The only difference between them is when they are quoted in quotes—see manual page for bash

$? is exit status of last command

$$ is the process ID of the current shell

Example shell script:

#! /bin/sh
echo $0 is the full name of this shell script
echo first parameter is $1
echo first parameter is $2
echo first parameter is $3
echo total number of parameters is $#
echo process ID is $$

Variables: use Braces ${...}

It’s good to put braces round a variable name when getting its value

Then no problem to join its value with other text:

$ test=123
$ echo ${test}
123
# No good, variable $test456 is undefined:
$ echo $test456
$ echo ${test}456
123456

Braces and Parameters after $9

Need braces to access parameters after $9:

$ cat paramten
#! /bin/sh
echo $10
echo ${10}
$ ./paramten a b c d e f g h i j
a0
j

Notice that $10 is the same as ${1}0, i.e., the first parameter “a” then the literal character zero “0”

More about Quoting

Double quotes: "..." stops the special behaviour of all special characters, except for:

variable interpretation ($) backticks (`) — see slide 24 the backslash (\)

Single quotes '...':

stop the special behaviour of all special characters Backslash:

preserves literal behaviour of character, except for newline Putting “\” at the end of the line lets you continue a long line on more than one physical line, but the shell will treat it as if it were all on one line.

OSSI — Shell Programming — ver. 1.9 – p. 23/85

Command Substitution

Enclose command in $(...) or backticks:`...`

Means, “Execute the command called within these quotes and put the output back here.”

Here is an example using expr:

$ expr 3 + 2
5

$ i=expr 3 + 2 # error: try execute command ‘3’

$ i=$(expr 3 + 2) # correct
$ i=`expr 3 + 2`  # also correct

Command Substitution—Another Example

$ hostname
nickpc.tyict.vtc.edu.hk
$ h=hostname
$ echo $h
hostname
$ h=$(hostname)
$ echo $h
nickpc.tyict.vtc.edu.hk

Conditions—String Comparisons

All programming languages depend on conditions for if statements and for while loops

Shell programming uses a built-in command which is

either test or [...]

Examples of string comparisons:

[ "$USER" = root ]     # true if the value of $USER is "root"
[ "$USER" != root ]   # true if the value of $USER is not equal to "root"
[ -z "$USER" ]         # true if the string "$USER" has zero length
[ string1 < string2 ] # true if string1 sorts less than string2
[ string1 > string2 ] # true if string1 sorts greater than string2

Note: the spaces after the “[“ and before the “]” are essential.

Conditions—Integer Comparisons

Examples of numeric integer comparisons:

[ "$x" -eq 5 ]  # true if the value of $x is 5
[ "$x" -ne 5 ]  # true if the value of $x is not 5
[ "$x" -lt 5 ]  # true if $x is less than 5
[ "$x" -gt 5 ]  # true if $x is greater than 5
[ "$x" -le 5 ]  # true if $x is less than or equal to 5
[ "$x" -ge 5 ]  # true if $x is greater than or equal to 5

Note again that the spaces after the “[“ and before the “]” are essential.

Conditions—File Tests, NOT Operator

The shell provides many tests of information about files.

Do man test to see the complete list.

Some examples:

$ [ -f file ]   # true if file is an ordinary file
$ [ ! -f file ] # true if file is NOT an ordinary file
$ [ -d file ]   # true if file is a directory
$ [ -u file ]   # true if file esists and has SUID permission
$ [ -g file ]   # true if file esists and has SGID permission
$ [ -x file ]   # true if file esists and is executable
$ [ -r file ]   # true if file esists and is readable
$ [ -w file ]   # true if file esists and is writeable
$ [ file1 -nt file2 ] # true if file1 is newer than file2

Conditions—Combining Comparisons

Examples of combining comparisons with AND: -a and OR: -o, and grouping with \(...\)

# true if the value of $x is 5 AND $USER is not equal to root:
[ "$x" -eq 5 -a "$USER" != root ]
# true if the value of $x is 5 OR $USER is not equal to root:
[ "$x" -eq 5 -o "$USER" != root ]
# true if ( the value of $x is 5 OR $USER is not equal to root ) AND
# ( $y > 7 OR $HOME has the value happy )
[ \( "$x" -eq 5 -o "$USER" != root \) -a \
\( "$y" -gt 7 -o "$HOME" = happy \) ]

Note again that the spaces after the “[“ and before the “]” are essential.

Do man test to see the information about all the operators.

Arithmetic Assignments

Can do with the external program expr

… but expr is not so easy to use, although it is very standard and portable

Easier is to use the built in let command Examples:

$ let x=1+4
$ let x='1 + 4'
$ let 'x = 1 + 4'
$ let x="(2 + 3) * 5"     # now x is 25
$ let "x = 2 + 3 * 5"     # now x is 17
$ let "x += 5"             # now x is 22
$ let "x = x + 5"         # now x is 27; note do not need $ on rhs

Notice that you do not need to quote the special characters with let.

Quote if you want to use white space.

Arithmetic Expressions with $((...))

The shell interprets anything inside $((...)) as an arithmetic expression

You could calculate the number of days left in the year like this:

$ echo "There are \
$(( (365-$(date +%j)) / 7 )) weeks \
left till December 31"

You don’t need to put a dollar sign in front of variables in arithmetic expressions, but it is okay to put them there.

Arithmetic Conditions with ((...))

A (less portable) alternative to the arithmetic conditions in slide 27 is putting the expression in ((...))

So you can do

(( (3>2) && (4<=1) ))

instead of

[ \( 3 -gt 2 \) -a \( 4 -le 1 \) ]

Operators that work with let, $((...)) and ((...)) are:

+ - * / % << >> & | ~ ! ^
< > <= >= == !=

which have exactly the same effect as in the C programming language.

if Statement

Syntax:

if test-commands
then
     statements-if-test-commands-1-true
elif test-commands-2
then
     statements-if-test-commands-2-true
else
     statements-if-all-test-commands-false
fi

Example:

if grep nick /etc/passwd > /dev/null 2>&1
then
     echo Nick has a local account here
else
     echo Nick has no local account here
fi

while Statement

Syntax:

while test-commands
do
     loop-body-statements
done

Example:

i=0
while [ "$i" -lt 10 ]
do
     echo -n "$i "        # -n suppresses newline.
     i=`expr $i + 1`      # let "i = i + 1" also works.
done

for Statement

Syntax:

for name in words
do
     loop-body-statements
done

Example:

for planet in Mercury Venus Earth Mars Jupiter Saturn Uranus Neptune Pluto
do
     echo $planet
done

for Loops: Another Example

Here the shell turns *.txt into a list of file names ending in “.txt”:

for i in *.txt
do
     echo $i
     grep 'lost treasure' $i
done

You can leave the in words out; in that case, name is set to each parameter in turn:

i=0
for parameter
do
     let 'i = i + 1'
     echo "parameter $i is $parameter"
done

for Loops: second, C-like syntax

There is a second (less frequently used, and less portable) C-like for loop syntax:

for (( expr1 ; expr2 ; expr3 ))
do
     loop-body-statements
done

Rules: same as for arithmetic conditions—see slide 32

Example:

for (( i = 0; i < 10; ++i ))
do
     echo $i
done

    break and continue

Use inside a loop Work like they do in C break terminates the innermost loop; execution goes on after the loop continue will skip the rest of the body of the loop, and resume execution on the next itteration of the loop.

Blocks: {...}

A subshell is one way of grouping commands together, but it starts a new process, and any variable changes are localised

An alternative is to group commands into a block, enclosing a set of commands in braces: {...}

Useful for grouping commands for file input or output

See next slide for another application.

Error Handling: ||, && and exit

Suppose we want the user to provide exactly two parameters, and exit otherwise

A common method of handling this is something like:

[ $# -eq 2 ] || { echo "Need two parameters"; exit 1; }

Read this as “the number of parameters is two OR exit”

Works because this logical OR uses short-circuit Boolean evaluation; the second statement is executed only if the first fails (is false)

Logical AND “&&” can be used in the same way; the second statement will be executed only if the first is successful (true)

A note about blocks: must have semicolon “;” or newline at end of last statement before closing brace

  Output: echo and printf

To perform output, use echo, or for more formatting,

printf.

Use echo -n to print no newline at end. Just echo by itself prints a newline printf works the same as in the C programming language, except no parentheses or commas:

$ printf "%16s\t%8d\n" $my_string $my_number

Do man printf (or look it up in the bash manual page) to read all about it.

Input: the read Command

For input, use the built-in shell command read read reads standard input and puts the result into one or more variables If use one variable, variable holds the whole line Syntax:

read var1 var2...

Often used with a while loop like this:

while read var1 var2
do
     # do something with var1 and var2
done

Loop terminates when reach end of file

set: Splitting a Multi-Word Variable

Sometimes may want to split a multi-word variable into single-word variables

read won’t work like this:

MY_FILE_INFO=$(ls -lR | grep $file)
# ...
echo $MY_FILE_INFO | read perms \
links user group month day time filename

Use the builtin command set instead:

MY_FILE_INFO=$(ls -lR | grep $file)
# ...
set $MY_FILE_INFO
read=$1 perms=$2 links=$3 user=$4 group=$5
month=$6 day=$7 time=$8 filename=$9

More about set, and IFS

set splits its arguments into pieces (usually) at whitespace

It sets the first value as $1, the second as $2, and so on.

Note that you can change how set and the shell splits things up by changing the value of a special variable called IFS

IFS stands for Internal Field Separator

Normally the value of IFS is the string

“⟨space⟩⟨tab⟩⟨newline⟩”

Next slide shows how changing IFS to a colon let us easily split the PATH into separate directories: ⇝ next slide

Example: Changing IFS

Notice that here, I make the change to IFS in a subshell. I have simply typed the loop at the prompt. As I said in slide 17, changes in a subshell are local to the subshell:

 $ echo $PATH
/usr/bin:/bin:/usr/X11R6/bin:/home/nicku/bin
$ (IFS=:
> for dir in $PATH
> do
>      echo $dir
> done
> )
/usr/bin
/bin
/usr/X11R6/bin
/home/nicku/bin

case Statement

Similar to the switch statement in C, but more useful and more general

Uses pattern matching against a string to decide on an action to take

Syntax:

case expression in
     pattern1 )
         statements ;;
     pattern2 )
         statements ;;
     ...
esac

case Statement: Example

This example code runs the appropriate program on a graphics file, depending on the file extension, to convert the file to another format:

case $filename in
*.tif)
tifftopnm $filename > $ppmfile
;;
*.jpg)
tjpeg $filename > $ppmfile
;;
*)
echo -n "Sorry, cannot handle this "
echo "graphics format"
;;
esac

shift: Move all Parameters Up One Place

Sometimes we want to process command-line parameters in a loop

The shift statement is made for this

Say that we have four parameters: $1 has the value one, $2 has the value two, $3 has the value three and $4 has the value four

Then after executing the shift statement, the values are now: $1 has the value two, $2 has the value three and $3 has the value four, and $4 no longer exists.

shift: Move Parameters Up Many Places

You can give a number argument to shift: If before, we have four parameters: $1 has the value one, $2 has the value two, $3 has the value three and $4 has the value four After executing the statement:

$ shift 2

we have two parameters left: $1 has the value three and $2 has the value four; $3 and $4 no longer exist.

Command-Line Options—1

Sometimes we want to modify the behaviour of a shell script

For example, want an option to show more information on request could use an option “-v” (for “verbose”) to tell the shell script that we want it to tell us more information about what it is doing If script is called showme, then we could use our -v option like this:

$ showme -v

the script then shows more information.

For example, We might provide an option to give a starting point for a script to search for suid programs Could make the option -d directory If script is called findsuid, could call it like this:

$ findsuid -d /usr

to tell the script to start searching in the directory /usr instead of the current directory We could do this using shift, a while loop, and a case statement, like this:

while [ -n "$(echo $1 | grep '-')" ]
do
case $1 in
-v) VERBOSE=1 ;;
-d)
shift
DIRECTORY=$1
;;
*)  echo "usage: $0 [-v] [-d dir]"
exit 1 ;;
esac
shift
done

OSSI — Shell Programming — ver. 1.9 – p. 52/85

getopts: Command-Line Options—4

Problems with above solution: inflexibility: Does not allow options to be “bundled” together like

-abc instead of -a -b -c

Requires a space between option and its argument, i.e., doesn’t let you do -d/etc as well as -d /etc Better method: use the built-in command getopts:

while getopts ":vd:" opt
do
     case opt in
         v) VERBOSE=1 ;;
         d) DIRECTORY=$OPTARG ;;
         *)  echo "usage: $0 [-v] [-d dir]"
              exit 1 ;;
     esac
done
shift $((OPTIND - 1))

getopts: Command-Line Options—5

getopts takes two arguments:

first comes the string that can contain letters and colons.

Each letter represents one option A colon comes after a letter to indicate that option takes an arguement, like -d directory A colon at the beginning makes getopts less noisy, so you can provide your own error message, as shown in the example.

The second is a variable that will hold the option (without the hyphen “-”)

Shift out all processed options using the variable OPTIND, leaving any other arguments accessible

Search for getopts in the bash man page

Temporary Files: mktemp

Sometimes it is convenient to store temporary data in a temporary file

The mktemp program is designed for this

We use it something like this:

TMPFILE=$(mktemp /tmp/temp.XXXXXX) || exit 1

mktemp will create a new file, replacing the “XXXXXX” with a random string

Do man mktemp for the complete manual.

Signals that may Terminate your Script

Many key strokes will send a signal to a process

Examples:

[figure: Control-C]sends a SIGINT signal to the current process running in the foreground

[figure: Control-\]sends a SIGQUIT signal

When you log out, all your processes are sent a SIGHUP (hangup) signal

If your script is connected to another process that terminates unexpectedly, it will receive a SIGPIPE signal

If anyone terminates the program with the kill program, the default signal is SIGTERM

Signals: trap

Sometimes you want your script to clean up after itself nicely, and remove temporary files Do this using trap

Signals: trap Example

Supose your script creates some temporary files, and you want to remove them if your script recieves any of these signals

You can “catch” the signal, and remove the files when the signals are received before the program terminates

Suppose the temporary files have names stored in the variables TEMP1 and TEMP2

Then you would trap these signals like this:

trap "rm $TEMP1 $TEMP2" HUP INT QUIT PIPE TERM

Conveniently, (but not very portably), bash provides a “pretend” signal called EXIT; can add this to the list of signals you trap, so that the temporary files will be removed when the program exits normally.

Functions

The shell supports functions and function calls

A function works like an external command, except that it does not start another process

Syntax:

function functname

{

     shell commands

} Or:

functname ()

{

     shell commands

}

Parameters in Functions

Work the same as parameters to entire shell script First parameter is $1, second is $2,…, the tenth parameter is ${10}, and so on. $# is the number of parameters passed to the function As with command line parameters, they are read-only Assign to meaningful names to make your program more understandable

Example, Calling a Function

This is a simple example program:

#! /bin/sh
function cube
{
     echo $(($1 * $1 * $1))
}
j=$(cube 5)
echo $j      # Output is 125

Note the use of command substitution to get a return value.

The function prints result to standard output.

Debugging Shell Scripts—1

If you run the script with:

$ sh -v script

then each statement will be printed as it is executed If you run the script with:

$ sh -x script

then an execution trace will show the value of all variables as the script executes.

Debugging Shell Scripts—2

Use echo to display the value of variables as the program executes You can turn the -x shell option on in any part of your script with the line:

set -x

and turn it off with:

set +x

The book Learning the bash Shell includes a bash shell debugger if you get desperate

Writing Shell Scripts

Build your shell script incrementally:

Open the editor in one window, have a terminal window open in which to run your program as you write it Test as you implement: this makes shell script development easy Do not write a very complex script, and then begin testing it!

Use the standard software engineering practice you know:

Use meaningful variable names, function names Make your program self-documenting Add comment blocks to explain obscure or difficult parts of your program

OSSI — Shell Programming — ver. 1.9 – p. 64/85

Useful External Programs—1

Each of these has a manual page, and many have info manuals. Read their online documentation for more information.

awk — powerful tool for processing columns of data basename — remove directory and (optionally) extension from file name cat — copy to standard output cut — process columns of data du — show disk space used by directories and files egrep, grep — find lines containing patterns in files find — find files using many criteria

Useful External Programs—2

last — show the last time a user was logged in lastb — show last bad log in attempt by a user rpm — rpm package manager: manage software package database sed — stream editor: edit files automatically sort — sort lines of files by many different criteria tr — translate one set of characters to another set uniq — replace repeated lines with just one line, optionally with a count of the number of repeated lines

Regular Expressions

Many programs and programming languages use regular expressions, including Java 1.4, Perl and VB.NET (plus many others; even MS Word uses regular expressions under Find → Advanced)

These programs use regular expressions:

   grep, egrep, sed, awk

Regular expressions provide a powerful language for manipulating data and extracting important information from masses of data

What is in a Regular Expression?

There are two types of character in a regular expression:

Metacharacters These include:

    *\. +? ^()[{|

Ordinary, literal characters: all the other characters

Literal characters

Find all lines containing "chan" in the password file:

$ grep chan /etc/passwd

The regular expressions is "chan" It is made entirely of literal characters It matches only lines that contain the exact string It will match lines containing the words chan, changed,

merchant, mechanism,. . .

Character Classes: [...]

A character class represents one character

Examples:

# Find all words in the dictionary that contain a vowel:
$ grep "[aeiou]" /usr/share/dict/words
# Find all lines that contain a digit:
$ grep "[0123456789]" /usr/share/dict/words
# Find all lines that contain a digit:
$ grep "[0-9]" /usr/share/dict/words
# Find all lines that contain a capital letter:
$ grep "[A-Z]" /usr/share/dict/words

Negated Character Classes: [^...]

Examples of negated character classes:

# Find all words in the dictionary
# that contain a character that is not a vowel:
$ grep "[^aeiou]" /usr/share/dict/words
# Two ways of finding all lines that contain
# a character that is not a digit:
$ grep "[^0123456789]" /usr/share/dict/words
$ grep "[^0-9]" /usr/share/dict/words
# Find all lines that contain a character
# that is not a digit, or a letter
$ grep "[^0-9a-zA-Z]" /usr/share/dict/words

Remember: each set of square brackets represents exactly one character.

Match Any Character

The dot “.” matches any single character, except a newline.

The pattern ’.....’ matches all lines that contain at least five characters

Matching the Beginning or End of Line

To match a line that contains exacly five characters:

$ grep '^.....$' /usr/share/dict/words

The hat, ^represents the position right at the start of the line The dollar $ represents the position right at the end of the line. Niether ^nor $ represents a character They represent a position Sometimes called anchors, since they anchor the other characters to a specific part of the string

Matching Repetitions: *, ?, +, {n}, {n,m}

To match zero or more: a* represents zero or more of the lower case letter a, so the pattern will match "" (the empty string), “a”,

“aa”, “aaaaaaaaaaaaaaa”, “qwewtrryu” or the

“nothing” in front of any string! To match one or more: ‘a+’ matches one or more “a”s ‘a?’ matches zero or one “a” ‘a{10}’ matches exactly 10 “a”s ‘a{5,10}’ matches between 5 and 10 (inclusive) “a”s

Matching Alternatives: “|”

the vertical bar represents alternatives: The regular expresssion ‘nick|albert|alex’ will match either the string “nick” or the string “albert” or the string “alex” Note that the vertical bar has very low precedence: the pattern ‘^fred|nurk’ matches “fred” only if it occurs at the start of the line, while it will match “nurk” at any position in the line

Putting it All Together: Examples

Find all words that contain at least two vowels:

$ grep '[aeiou].*[aeiou]' /usr/share/dict/words

Find all words that contain exactly two vowels:

$ egrep \
'[^aeiou]*[aeiou][^aeiou]*[aeiou][^aeiou]*' \
/usr/share/dict/words

Find all lines that are empty, or contain only white space:

$ grep '^\s*$' file

Basic awk

awk is a complete programming language

Mostly used for one-line solutions to problems of extracting columns of data from text, and processing it

A complete book is available on awk at

http://www.ssc.com/ssc/eap/ .

What Does awk Do?

awk reads file(s) or standard input one line at a time, and automatically splits the line into fields, and calls them

$1, $2,. . . , $NF

NF is equal to the number of fields the line was split into $0 contains the whole line awk has an option -F that allows you to select another pattern as the field separator

Normally awk splits columns by white space To execute code after all lines are processed, create an END block.

awk Examples

Print the sizes of all files in current directory:

 ls -l | awk '{print $5}'

Add the sizes of all files in current directory:

 ls -l | awk '{sum += $5} END{print sum}'

Print only the permissions, user, group and file names of files in current directory:

 ls -l | awk '{print $1, $3, $4, $NF}'

sed—the Stream Editor

sed provides many facilities for editing files The substitute command, s///, is the most important The syntax (using sed as an editor of stanard input), is:

$ sed 's/original/replacement/'

Example: replace the first instance of Windows with Linux on each line of the input:

 sed 's/Windows/Linux/'

Example: replace all instances of Windows with Linux on each line of the input:

 sed 's/Windows/Linux/g'

sed—Backreferencees

You can match part of the original in a sed substitute command, and put that part back into the replacement part.

You enclose the part you want to refer to later in

\(...\)

You can get the first value in the replacement part by \1, the second opening parenthesis of \(...\) by \2, and so on.

sed—Backreferencees: Example

If you do find /etc | xargs file -b, you will get

a lot of output like this:

symbolic link to bg5ps.conf.zh_TW.Big5
symbolic link to rc.d/rc.local
symbolic link to rc.d/rc
symbolic link to rc.d/rc.sysinit
symbolic link to ../../X11/xdm/Xservers

If you want to edit each line to remove everything after “symbolic link”, then you could pipe the data through sed like this:

$ find /etc | xargs file -b \
| sed 's/\(symbolic link\).*/\1/'

See slide 83 for an application

find Examples

Count the number of unique manual pages on the computer:

$ find /usr/share/man -type f | wc -l

Print a table of types of file under the /etc directory, with the most common file type down at the bottom:

$ find /etc | xargs file -b \
  | sed 's/\(symbolic link\).*/\1/' \
  | sort \
  | uniq -c \
  | sort -n

A find Example with Many Options

Set all directories to have the access mode 771, set all backup files (*.BAK) to mode 600, all shell scripts (*.sh) to mode 755, and all text files (*.txt) to mode 644:

$ find . \( -type d        -a exec chmod 771 {} \; \) -o \
\( -name "*.BAK" -a exec chmod 600 {} \; \) -o \
\( -name "*.sh"  -a exec chmod 755 {} \; \) -o \
\( -name "*.txt" -a exec chmod 644 {} \; \)

rpm Database Query Commands

The rpm software package management system includes a database with very detailed information about every file of every software package that is installed on the computer.

You can query this database using the rpm command.

The manual page does not give the complete picture, but there is a book called Maximum RPM that comes on the Red Hat documentation cd

This package is installed on ictlab

You can see the appropriate section at this url:

http://nicku.org/doc/maximum-rpm-1.0/html/s1-rpm-query-parts.html